OpenStudy (anonymous):

can someone factor 3x^2 - 11x + 6

OpenStudy (anonymous):

(3x+a)(x+b) a*b =6 3b +a =-11 find a and b

hero (hero):

Split the middle term 3x^2 - 9x - 2x + 6 3x(x-9) - 2(x-9) (x-9)(3x-2)

hero (hero):

Algebraic, it's ac = 18 a + c = -11

hero (hero):

@algebraic!

OpenStudy (anonymous):

thankss @Hero ! :)